# Traditional Methods Outperform Generative LLMs at Forecasting Credit Ratings

### v1

@@ 3.1.4. Macroeconomic data (M)
Adverse events in the world economy can also impact a company’s ability to repay its debt. Many external forces can affect a company’s future creditworthiness, however, we have identified three key areas from prior research in the area (Carling et al., 2007; Taylor et al., 2021): labour statistics, interest rates and foreign exchange data. For the labour statistics, we used the Bureau of Labour Statistics dataset [^6] Labour Statistics: https://tinyurl.com/y94d52xk. For the interest rate and foreign exchange data, we used the Federal Reserve Bank Reports [^7] Interest Rate Data: ${}^{\text{,}}$ [^8] Foreign Exchange Data: https://tinyurl.com/a38rmzd8.
3.2. Dataset Construction
@@
To ensure consistency with the periodicity of SEC filings, all data in our study is aligned every quarter. This dataset extends from Q1 1994, marking the beginning of the SEC’s electronic processing of filings, to Q2 2017, the last point at which we have credit rating data from Compustat Capital IQ. Not every company in our dataset has a complete and continuous record of data. Consequently, we have excluded instances with incomplete data. A significant outcome of this approach is that when the number of lagged quarters used in the task is increased, the number of valid samples diminishes. This reduction is due to the lower probability of having complete data across many consecutive quarters, compared to when only the most recent quarter is considered.
@@
The class imbalance in credit rating data is highly skewed. 93.4% of the companies in our dataset maintain the same score. Whilst it is common convention to use oversampling techniques such as SMOTE (Chawla et al., 2002) to aid credit rating prediction models (Pamuk and Schumann, 2023; Wang, 2022; Zhao et al., 2024). There is no consensus on how SMOTE could be applied to high-dimensional text embeddings. As a result, we balanced the classes for this task, reducing the size of the datasets. The train test splits run from 01-01-1994 to 31-12-2012, the validation splits run from 01-01-2013 to 31-12-2014, and the test splits from 01-01-2015 to 31-12-2016.
@@
The MDA section of an SEC Filing, despite only constituting a small part of the filing, is still very long. The average MDA section considered in our task is 13,267 tokens long using a BPE tokenizer (Sennrich et al., 2016). As a result, when a model could not encode all of the tokens in the MDA section, only the first part of the text was encoded, up to either a set length or the maximum input sequence length.
4. Methodology
@@
We deploy two frameworks to test different architectural methodologies on this task We maintain the same data available to each variation of the frameworks. The first framework is a feature-based discriminative approach that uses a more traditional boosting-tree model and tests the different ways to encode the textual data. The second uses generative language models and prompting to output one of a fixed list of labels through a greedy search algorithm (Appendix B).
4.1. Task Description
@@
The objective of the task is to predict the credit rating, $\hat{R}_{t}$, at time $t$. The function can be represented as follows:
@@
Here, $T_{t-i}$ represents the text data, $R_{t-i}$ represents the historical credit rating data, $N_{t-i}$ represents the numeric data - both fundamental and macroeconomic. $i$ varies from 1 to $p$, with $p$ indicating the number of past quarters considered (1 to 4 quarters in this study). Furthermore, $f$ is the predictive function to convert our input data into an estimate. An ablation study is conducted to evaluate the impact of different data types on the prediction accuracy. In this study, the function $f$ is tested under various configurations: using only text data $T_{t-i}$, using combinations of historical ratings $R_{t-j}$, and numeric data $N_{t-k}$. This approach helps to determine the relative importance of each type of data.
4.2. Boosting-Tree Baseline
@@
To test the abilities of generative language models, it is necessary to benchmark the performance against a relatively well-understood and robust algorithm. We select XGBoost (Chen and Guestrin, 2016), a model that has been widely adopted in many domains (Talukder et al., 2023; Dong et al., 2023; Joshi et al., 2024). The supervised model takes as input the normalised fundamental, macroeconomic and text data, and outputs the most likely label.
4.3. Text Encoders
Data type Av. Quarters
@@ 4.3. Text Encoders
Numeric M + F + CR 52.81 48.29 53.32 53.99 55.65
CR 44.68 41.92 43.57 46.48 46.73
All LM Lex 50.64 46.83 51.19 52.13 52.42
Emotion Lex 51.35 49.76 52.41 51.60 51.61
LDA 50.88 50.25 52.34 50.76 50.16
Clusters 53.61 50.73 54.64 54.13 55.96
BERT Emotion 52.83 48.23 52.92 55.35 54.83
BERT 52.69 50.51 54.09 52.29 53.89
Text only LM Lex 34.42 36.56 33.42 30.85 36.83
Emotion Lex 35.51 32.93 34.48 36.17 38.44
LDA 34.75 36.59 35.01 30.32 37.10
Clusters 38.12 39.76 38.01 38.86 35.83
BERT Emotion 36.01 36.83 35.81 36.17 35.22
BERT 32.65 32.20 28.91 36.17 33.33
[…]
4.3.1. Loughran McDonald Lexicon (LM Lex)
(Loughran and McDonald, 2011) - The LM Lexicon is widely recognized in finance. Given its prevalence, it is crucial to compare its effectiveness with more advanced methods. The lexicon classifies words into four domains: Positivity, Negativity, Litigiousness, and Uncertainty. However, the simple language modelling technique classifies phrases like ”The debt increased last quarter” as neutral. The LM text representation in this work is the count of how many words from each sentiment occur in the document, normalized by the maximum value in the train set.
4.3.2. NRC Lexicon (Emotion Lex)
(Mohammad, 2020) - The NRC lexicon is a general lexicon which scores words for ten different emotions and sentiments and is used for multi-domain problems. The text representation is created in the same way as the LM Lexicon.
4.3.3. Latent Dirichlet Allocation (LDA)
(Blei et al., 2003) A prominent topic modelling technique often employed to assign topics to sequences of text. Although the field of topic modelling has evolved, LDA continues to serve as a robust baseline for evaluating new topic models. To create the text representation we created a vector where each of the dimensions represented the probability that the text belonged to each one of 25 different topics.
4.3.4. High-density Embedding Clusters (Clusters)
(Grootendorst, 2022) - Contextual word embeddings have provided a good basis for topic modelling (Sia et al., 2020; Grootendorst, 2022). Sentence embeddings have also been used to discern domain type from text (Aharoni and Goldberg, 2020). (Drinkall et al., 2022) extended this work to generate features from clusters of sentence embeddings in a COVID-19 caseload prediction task. For this task, each sentence of each filing in the train set was encoded into embeddings space using a all-mpnet-base-v2 (Reimers and Gurevych, 2019), the dimensionality was then reduced using UMAP (McInnes et al., 2020), and the HDBSCAN clustering algorithm (Campello et al., 2013) was used for form 100 distinct clusters. Examples of the clusters will be included in supplementary material. Then each filing in the train, validation and test set was split into sentences and then transformed into the embedding space described above. The overall text representation was the average of the representations of each sentence, and the representation of each sentence was the probability distribution that the sentence belonged to each of the 100 clusters.
4.3.5. DistilRoBERTa Emotion (BERT Emotion)
(Hartmann, 2022) - In this work a DistilRoBERTa model has been fine-tuned on an emotion classification task that includes Ekman’s basic emotions (Ekman, 1992), plus the neutral label: anger, disgust, fear, joy, neutral, sadness, surprise. This methodology provides a more complex comparison of whether emotions are a useful feature base for predicting a change in credit ratings. The SEC Filings are then chunked into 512 token sequences and classified according to the probability that that chunk can be associated with each emotion. The average across all chunks is taken as the final text representation of each filing.
4.3.6. BERT
(Devlin et al., 2019) - The first 512 tokens of each SEC Filing are passed into BERT and the average embedding from the final layer is taken as the textual representation of that filing.
Model Data type Average Quarters Feature Importances
1 2 3 4 Macro Fund Text
XGBoost Numeric 52.81 48.29 53.32 53.99 55.65 0.6593 0.3407 -
All (Clusters) 53.86 50.73 54.64 54.13 55.96 0.1174 0.1880 0.6946
Text Only (Clusters) 38.12 39.76 38.01 38.86 35.83 - - -
Numeric Features + GPT4o Text Estimate 53.81 52.68 50.93 53.99 57.26 0.5877 0.3023 0.1010
XGBoost Numeric Estimate + GPT4o Text Estimate 51.75 53.66 51.19 50.80 51.34 0.5727 0.4273
GPT-4o Numeric 31.41 33.66 31.57 31.12 29.30 - - -
All 40.24 43.90 40.32 38.83 37.90 - - -
Text Only 49.64 49.27 52.23 52.39 44.62 - - -
Text + XGBoost Numeric Estimate 32.27 33.90 30.77 36.70 27.69 - - -
Union of XGBoost-Numeric & GPT-4o Text-Only 69.87 70.49 73.21 71.54 64.25 - - -
[…]
4.4. Generative Framework
Generative LLMs have seen significant recent progress. Given this development, it is important to see if these models can discern whether there is likely to be a change in the perceived risk of a company, and if so, what is the best methodology for achieving high performance. This section is methodologically distinct from the other text encoding methods described above since the numerical data is encoded into text before the model attempts to solve the task through the use of prompting. The prompts used in the following section are included in the Appendix F, and follow the best practice from existing literature (OpenAI, 2024; Lin et al., 2024).
Whilst LLMs perform very well in 0-shot settings (Kojima et al., 2023), there is significant evidence that shows that LLMs perform better in a k-shot setting (Clark et al., 2018); the ARC benchmark uses 25-shot prompts in the Eleuther AI evaluation harness (Gao et al., 2023). The problem with deploying a k-shot framework in this setting is that the SEC Filings are very long (13,267 tokens). Despite the increase in the context-window length of some newer LLMs, many new models are capped at 8192 tokens or below (Wang and Komatsuzaki, 2021; Jiang et al., 2023; Touvron et al., 2023; Singh et al., 2024), and some other studies have shown performance deterioration as the input sequence increases (Li et al., 2024). Fitting several examples of the task in the input sequence is impossible for many of the data samples, which means that k-shot performance is not reported for this task.
We tested the following models using the prompting structure laid out in Appendix F: gpt-3.5-turbo-0125, gpt-4-turbo-2024-04-09 and gpt-4o-2024-05-13, Llama-3 8B. These models provide a good representation of the current state-of-the-art (Chiang et al., 2024).
4.4.1. LoRA Adaptation
To adapt the LLMs to the task we use LoRA (Low Rank Adaptation) (Hu et al., 2021) fine-tuning. This technique involves optimizing the rank-decomposition matrices, $A$ $B$, of the change in model weights ($\Delta W$), where $W^{{}^{\prime}}$ are the new model weights and $W$ are the pre-trained weights.
(1) $W^{{}^{\prime}}$ $=W+\Delta W$
(2) $=W+BA$
@@ 4.4.1. LoRA Adaptation
Num. Llama-3 8B 32.31 35.12 35.81 28.99 29.30
Llama-3 8B LoRA 35.51 35.37 34.22 37.77 34.68
GPT-3.5 32.63 32.93 33.42 31.91 32.26
GPT-4 34.08 34.15 32.89 32.98 36.29
GPT-4o 31.41 33.66 31.57 31.12 29.30
All Llama-3 8B 35.50 35.37 35.81 36.97 33.87
Llama-3 8B LoRA 37.47 36.39 37.01 37.64 38.83
GPT-3.5 44.52 49.02 42.44 46.01 40.59
GPT-4 38.28 39.76 36.07 38.03 39.25
GPT-4o 40.24 43.90 40.32 38.83 37.90
Text Llama-3 8B 35.57 35.37 36.07 36.97 33.87
Llama-3 8B LoRA 36.96 38.10 36.75 37.11 35.87
GPT-3.5 46.36 47.32 47.48 45.48 45.16
GPT-4 48.48 47.80 48.54 50.27 48.12
GPT-4o 49.64 49.27 52.23 52.39 44.62
[…]
5. Results
@@
The results from the XGBoost baseline are outlined in Table 1. Firstly, it is clear that there is some information in the text since almost all text encoding methods perform above chance. None of the individual text feature sets outperform the numeric baselines, showing that the fundamental and macroeconomic variables are more important to the prediction than the text. However, when all of the features are combined we see a performance increase when the high-density cluster features are used in conjunction with the numerical information.
@@
The generative models show some interesting behaviour in Table 3. Using a 0-shot prompt, the performance using numerical information is near random. Indeed the performance is higher on the GPT-class models when only text is used as opposed to when all data types are used. This shows that the numeric information is harming the predictive performance. Interestingly, GPT-3.5 - an older model - performs the best when all data is considered. It also appears that LoRA enables slightly better relative performance on understanding numerical data - Llama-3 8B LoRA is the best-performing model on entirely numerical information and is the only model with no performance degradation when all features are considered as opposed to just text. However, the main conclusion from Table 3 is that generative models are very good relatively at decoding text data for this task.
@@
Table 2 takes the best-performing text features from the XGBoost framework, Clusters, and compares the model to the best-performing generative model, GPT-4o. Interestingly, GPT-4o utilises the text alone much better than any of the other encoder-based methods, however when all features are considered the XGBoost-Clusters configuration is the best-performing methodology. In addition to that, the models are also picking up on different signals. The proportion of samples where at least one of the XGBoost-Numeric and GPT-4o Text Only is correct (69.87) is significantly higher than any of the individual models. As a result, we provide comparisons where the estimate and probability of the XGBoost-Numeric and GPT-4o Text-Only are included in the prompt or feature set. Both the estimate and the probability that each model assigns to the estimate are used in the prompt or as features. None of the estimate combination methods outperform the XGBoost-Clusters framework, but it seems reasonable to suggest that future ensemble methods may be able to capture some of this performance gap.
6. Interpretability
One of the major disadvantages of using generative LLMs is that, to a large degree, they are black boxes. While some work has successfully used attention weights to analyse generated prompts (Serrano and Smith, 2019; Moradi et al., 2019), many of the best-performing LLMs are only accessible via APIs (Chiang et al., 2024), which remove the ability to analyse LLMs’ attention weights. Boosting trees, using known and interpretable features, on the other hand, provide increased interpretability. Regulation in major economies is moving towards a system that emphasizes explainability almost as much as absolute performance (Commission, 2020; US Congress, 2022; UK Secretary of State for Digital, Culture, Media and Sport, 2022); methods that provide a window into the statistical associations that the model infers during its decision-making process are essential. As well as providing the best performance of all the model configurations in this paper, the XGBoost-Clusters framework provides end-users with the ability to understand and explain the decisions that made. This section provides an example of how we can use this framework to understand the reasons behind decisions.
(a) PDP of ”rating_sp_ratings_moodys” cluster & ”Up” class.
(b) PDP of ”receivables_percent_owned_offs” cluster & ”Down” class.
Figure 2. Partial Dependence Plots (PDP) of text-based features against different target classes.

### v2

@@ 3.0.4 Macroeconomic data (M)
Adverse events in the world economy can also impact a company’s ability to repay its debt. Many external forces can affect a company’s future creditworthiness, however, we have identified three key areas from prior research in the area Carling et al. (2007); Taylor et al. (2021): labour statistics, interest rates and foreign exchange data. For the labour statistics, we used the Bureau of Labour Statistics dataset [^6] Labour Statistics: https://tinyurl.com/y94d52xk. For the interest rate and foreign exchange data, we used the Federal Reserve Bank Reports [^7] Interest Rate Data: ${}^{\text{,}}$ [^8] Foreign Exchange Data: https://tinyurl.com/a38rmzd8.
3.1 Dataset Construction
@@
To maintain consistent periodicity in SEC filings, all data is aligned quarterly. The dataset spans from Q1 1994, when the SEC began electronic processing of filings, to Q2 2017, the last period with credit rating data from Compustat Capital IQ. Companies with incomplete records were excluded. As a result, when the number of lagged quarters used in the task is increased, the number of valid samples diminishes. This reduction is due to the lower probability of having complete data across many consecutive quarters, compared to when only the most recent quarter is considered.
@@
Credit rating data is highly imbalanced, with 93.4% of companies maintaining the same score. While oversampling techniques like SMOTE Chawla et al. (2002) are common for credit rating prediction Pamuk and Schumann (2023); Wang (2022); Zhao et al. (2024), their application to text embeddings lacks consensus. To address this, we balanced the classes, reducing the dataset size. Training data spans Q1 1994 to Q4 2012, validation from Q1 2013 to Q4 2014, and testing from Q1 2015 to Q4 2016. The dataset was made from 23 years and the size is representative of many other tasks in NLP (Table 4).
@@
The MDA section of an SEC Filing, despite only constituting a small part of the filing, is still very long. The average MDA section in our task is 13,267 tokens long using a BPE tokenizer Sennrich et al. (2016). As such, when a model was not able to encode all of the tokens, only the first part of the text was encoded.
4 Methodology
@@
We deploy two frameworks to test different architectural methodologies on this task. The same data are provided to each of the frameworks. The first framework is a feature-based discriminative approach that uses a more traditional boosting-tree model and tests the different ways to encode the textual data. The second uses generative LLMs and prompting to output one of a fixed list of labels through a greedy search algorithm (App. A).
4.1 Task Description
@@
The objective is to predict the credit rating, $\hat{R}_{t}$, at time $t$. The function can be represented as follows:
@@
Here, $T_{t-i}$ represents the text data, $R_{t-i}$ represents the historical credit rating data, $N_{t-i}$ represents the numeric data - both fundamental and macroeconomic. $i$ varies from 1 to $p$, with $p$ indicating the number of past quarters considered (1 to 4 quarters in this study). Furthermore, $f$ is the predictive function to convert the input data into an estimate. An ablation study is conducted to evaluate the impact of different data types on the prediction accuracy. In this study, the function $f$ is tested under various configurations: using only text data $T_{t-i}$, using combinations of historical ratings $R_{t-j}$, and numeric data $N_{t-k}$. This approach helps to determine the relative importance of each type of data.
4.2 Boosting-Tree Baseline
@@
To test the abilities of generative LLMs, it is necessary to benchmark the performance against a relatively well-understood and robust algorithm. We select XGBoost Chen and Guestrin (2016), a model that has been widely adopted in many domains Talukder et al. (2023); Dong et al. (2023); Joshi et al. (2024). The supervised model takes as input the normalized fundamental, macroeconomic and text data, and outputs the most likely label. We describe other more complex neural network architectures that failed to learn this task in Appendix D. Due to the restricted dataset size, the models outlined in the Appendix were unable to learn the task before overfitting the training dataset.
4.3 Text Encoders
To test and identify which of the traditional encoders performs best we trialled a series of standard methodologies.
The Loughran McDonald Lexicon (LM) Loughran and McDonald (2011) is widely recognized in finance. Given its prevalence, it is crucial to compare its effectiveness with more advanced methods. The lexicon classifies words into four domains: Positivity, Negativity, Litigiousness, and Uncertainty. However, the simple language modelling technique classifies phrases like "The debt increased last quarter" as neutral. The LM text representation in this work is the document word count from each sentiment, normalized by the maximum value in the training set.
Latent Dirichlet Allocation (LDA) is a widely-used topic modeling method that identifies latent topics within text Blei et al. (2003). It operates by assuming that each document is a mixture of topics and that each topic is a distribution over words. Despite advancements in topic modeling, LDA remains a reliable baseline for evaluating newer models. In this paper, the features represent texts as probability distributions over 25 topics, with each dimension indicating the likelihood of the text belonging to a specific topic.
High-density Embedding Clusters (HEC) leverage the natural language understanding of LLMs but reduce the dimensionality of the input feature. HEC provides a good basis for topic modelling Sia et al. (2020); Grootendorst (2022). Sentence embeddings have also been used to discern domain type from text Aharoni and Goldberg (2020). Drinkall et al. 2022 extended this work to generate features from clusters of sentence embeddings in a COVID-19 caseload prediction task. For this task, each sentence of each filing in the training set was encoded into embeddings space using a all-mpnet-base-v2 Reimers and Gurevych (2019), the dimensionality was then reduced using UMAP McInnes et al. (2020), and the HDBSCAN clustering algorithm Campello et al. (2013) was used for form 100 distinct clusters. An example of the cluster features is displayed in Figure 1. Then each filing in the train, validation and test set was split into sentences and then transformed into the embedding space described above. The overall text representation was the average of the representations of each sentence, and the representation of each sentence was the probability distribution that the sentence belonged to each of the 100 clusters.
To understand the extent to which emotion scores play a significant role in forecasting the next credit rating. We used a DistilRoBERTa model ($\boldsymbol{E_{\text{DRoBERTa}}}$) that had been fine-tuned on an emotion classification task Hartmann (2022). The SEC Filings are then chunked into 512 token sequences and classified according to the probability that that chunk can be associated with each emotion. The average across all chunks is taken as the final text representation of each filing.
We also trialled a pooled MP-NET representation Song et al. (2020) by chunking the text into 512 token segments and averaging over the pooled representation of each chunk. In well-established benchmarks Muennighoff et al. (2023), MP-NET embeddings have performed strongly for their size and provide a baseline comparison to the HEC features derived from the MP-NET model.
4.4 Generative Framework
Given recent advancements in generative LLMs, we evaluate whether these models can identify changes in a company’s perceived risk and determine the best methodology for achieving high performance. This approach differs from other text encoding methods discussed earlier, as numerical data is converted into text format for the model to process using prompts. The prompts used in the following section are included in the Appendix E, and follow the best practice from existing literature OpenAI (2024); Lin et al. (2024). Sui et al. (2024) showed that contextual information about the tabular features enables a 0-shot framework to outperform 1-shot prompting methodology.
While LLMs perform very well in 0-shot settings Kojima et al. (2023), there is significant evidence that shows that LLMs perform better in a k-shot setting Clark et al. (2018); the ARC benchmark uses 25-shot prompts in the Eleuther AI evaluation harness Gao et al. (2023). The problem with deploying a k-shot framework in this setting is that the SEC Filings are very long (13,267 tokens). Despite the increase in the context-window length of some newer LLMs, many new models are capped at 8192 tokens or below Wang and Komatsuzaki (2021); Jiang et al. (2023); Touvron et al. (2023); Grattafiori et al. (2024), and some other studies have shown performance deterioration as the input sequence increases Li et al. (2024). Fitting several examples of the task in the input sequence is impossible for many of the data samples, which means that k-shot performance is not reported for this task.
We tested several models999gpt-3.5-turbo-0125, gpt-4-turbo-2024-04-09 and gpt-4o-2024-05-13, Llama-3 8B using the prompting structure laid out in Appendix E. The models provide a good representation of the current state-of-the-art Chiang et al. (2024).
4.4.1 LoRA Adaptation
To adapt the LLMs we use LoRA (Low Rank Adaptation) Hu et al. (2021) fine-tuning. This technique involves optimizing the rank-decomposition matrices, $A$ $B$, of the change in model weights ($\Delta W$), where $W^{{}^{\prime}}$ are the new model weights and $W$ are the pre-trained weights.
$W^{{}^{\prime}}$ $=W+\Delta W$ (1)
$=W+BA$ (2)
Data Features Av. Quarters
@@ 4.4.1 LoRA Adaptation
$\mathbf{N}$ ${M+F+C}$ 52.8 48.3 53.3 54.0 55.7
${C}$ 44.7 41.9 43.6 46.5 46.7
$\mathbf{A}$ LM 50.6 46.8 51.2 52.1 52.4
LDA 50.9 50.3 52.3 50.8 50.2
HEC 53.6 50.7 54.6 54.1 56.0
$E_{\text{DRoBERTa}}$ 52.8 48.2 52.9 55.4 54.8
MP-NET 51.0 46.8 52.5 56.1 48.4
$\mathbf{T}$ LM 34.4 36.6 33.4 30.9 36.8
LDA 34.8 36.6 35.0 30.3 37.1
HEC 38.1 39.8 38.0 38.9 35.8
$E_{\text{DRoBERTa}}$ 36.0 36.8 35.8 36.2 35.2
MP-NET 35.9 32.4 35.0 39.9 36.3
Table 1: The accuracy using the XGBoost model across different feature sets and text encoding methods. $\mathbf{N}$ refers to instances where only numeric information is used. $\mathbf{T}$ refers to text-based data types. $\mathbf{A}$ indicates all data types combined ($M+F+C+T$). Bold indicates the best results for each of the data configurations; underline indicates the best results across all configurations.
5 Results
Model Features Average Quarters Feature Importances
1 2 3 4 $\mathbf{M}$ $\mathbf{F}$ $\mathbf{T}$
XGBoost $\mathbf{N}$ 52.8 48.3 53.3 54.0 55.7 0.659 0.341 -
$\mathbf{A}_{\text{HEC}}$ 53.9 50.7 54.6 54.1 56.0 0.117 0.188 0.695
$\mathbf{T}_{\text{HEC}}$ 38.1 39.8 38.0 38.9 35.8 - - -
$\mathbf{N}+\mathbf{T}_{\text{GPT-4o Est.}}$ 53.8 52.7 50.9 54.0 57.3 0.588 0.302 0.101
$\mathbf{N}_{\text{XGB Est.}}+\mathbf{T}$ 51.8 53.7 51.2 50.8 51.3 0.573 0.427
GPT-4o $\mathbf{N}$ 31.4 33.7 31.6 31.1 29.3 - - -
$\mathbf{A}$ 40.2 43.9 40.3 38.8 37.9 - - -
$\mathbf{T}$ 49.6 49.3 52.2 52.4 44.6 - - -
$\mathbf{T}+\mathbf{N}_{\text{XGB Est.}}$ 32.3 33.9 30.8 36.7 27.7 - - -
XGBoost-$\mathbf{N}\cup\text{GPT-4o-}\mathbf{T}$ 69.9 70.5 73.2 71.5 64.3 - - -
Table 2: Accuracy across different model and data configurations. The notation is consistent to Table 1. Feature importances for $\mathbf{M},\mathbf{F},\mathbf{T}$ are impurity scores averaged across lags. $\mathbf{N}_{\text{XGB Est.}}$ and $\mathbf{T}_{\text{GPT-4o Est.}}$ represent the subscript model’s estimate and implied internal probability of that estimate using the features represented by the bold letter, both are the probability and estimate are used as features.
@@ 5 Results
$\mathbf{N}$ Llama 32.3 35.1 35.8 29.0 29.3
Llama-LoRA 35.5 35.4 34.2 37.8 34.7
GPT-3.5 32.6 32.9 33.4 31.9 32.3
GPT-4 34.1 34.2 32.9 33.0 36.3
GPT-4o 31.4 33.7 31.6 31.1 29.3
$\mathbf{A}$ Llama 35.5 35.4 35.8 37.0 33.9
Llama-LoRA 37.5 36.4 37.0 37.6 38.8
GPT-3.5 44.5 49.0 42.4 46.0 40.6
GPT-4 38.3 39.8 36.1 38.0 39.3
GPT-4o 40.2 43.9 40.3 38.8 37.9
$\mathbf{T}$ Llama 35.6 35.4 36.1 37.0 33.9
Llama-LoRA 37.0 38.1 36.8 37.1 35.9
GPT-3.5 46.4 47.3 47.5 45.5 45.2
GPT-4 48.5 47.8 48.5 50.3 48.1
GPT-4o 49.6 49.3 52.2 52.4 44.6
Table 3: Accuracy using the generative models. The notation is consistent to Table 1. All models are tested in 0-shot besides Llama-3 8B which is fine-tuned using LoRA.
@@
The results from the XGBoost baseline are outlined in Table 1. It is clear that there is some information in the text since almost all text encoding methods perform above chance. However, none of the individual text feature sets outperform the numeric baselines, indicating that fundamental and macroeconomic variables are more critical for prediction. Combining features yields a performance boost, particularly when HEC features are integrated with numerical data.
@@
Table 3 highlights intriguing behavior in generative models. With a zero-shot prompt, performance using only numerical data is near random. Interestingly, GPT-class models perform better using text alone than with all data types, suggesting that numerical information may hinder their predictive accuracy. GPT-3.5, despite being older, achieves the best performance with all features. It also appears that LoRA enables better relative performance on numerical data - Llama-3 8B LoRA is the best-performing model on entirely numerical information and is the only model with no performance degradation when all features are considered as opposed to just text. Overall, generative models excel at decoding text data for this task.
Table 2 takes the best-performing text features from the XGBoost framework, HEC, and provides a comparison to the best-performing generative model, GPT-4o. Interestingly, GPT-4o utilises the text alone much better than any of the encoder-based methods, but when all feature-types are considered the XGBoost-HEC configuration is the best-performing methodology.
@@
In addition, the models pick up on different signals. The final row of Table 2 shows that the proportion of samples where at least one of the XGBoost-$\mathbf{N}$ and GPT-4o-$\mathbf{T}$ is correct (69.87) is significantly higher than any of the individual models. As a result, we provide comparisons where the estimate and class probability of the XGBoost-$\mathbf{N}$ and GPT-4o-$\mathbf{T}$ are included in the prompt or feature set. Both the estimate and the probability that each model assigns to the estimate are used in the prompt or as features. The combination methods all underperform the XGBoost-$\mathbf{A}_{\text{HEC}}$ configuration, but this performance gap provides an opportunity for future research on ensembling methods.
6 Interpretability
One of the disadvantages of generative LLMs is that, to a large degree, they are black boxes. While some work has successfully used attention weights and internal model states to analyse generated prompts Serrano and Smith (2019); Wang et al. (2022), this is not a mature research area. Much of the mechanistic interpretability literature has focused on toy models Elhage et al. (2021), and while there has been some progress made on extracting features from larger models using sparse autoencoders Templeton et al. (2024), the field is still very far from completely solving interpretability within LLMs. In the absence of a complete solution, it is worth acknowledging the interpretable features that traditional methods use. Regulation in major economies increasingly emphasizes explainability alongside performance European Commission (2020); US Congress (2022); UK Secretary of State for DCMS (2022). The XGBoost-$\mathbf{A}_{\text{HEC}}$ framework not only achieves the best performance among the models in this paper but also enables users to interpret its decisions. This section provides an example of how we can use this framework to understand the reasons behind decisions.